# AI4Research: A Survey of Artificial Intelligence for Scientific Research

### v1

@@ Abstract:
${}^{\leavevmode{\mbox{=\hskip 0.0pt plus 0.70004pt}}}$ Corresponding Author
@@ 1 Introduction
In recent years, the rise of artificial intelligence (AI), particularly large language models (LLMs) like DeepSeek-R1 [263], has stimulated significant research in the field of reasoning. These breakthroughs have notably improved the models’ performance across diverse areas, including mathematical reasoning, programming, and interdisciplinary knowledge [724, 748, 616, 931, 947, 110]. Some of these models have even surpassed the Turing Test [352], marking a pivotal achievement in AI development. Inspired by these, a series of works attempts to explore advanced AI systems for innovative tasks, especially in the scientific discovery of new research [863, 887, 847, 948]. Earlier, the AI Scientist [507] introduces the concept of a fully automatic AI for research system, which divides the research process into three key stages: idea mining, experiment conduction, and academic writing. Initially, the system generates and evaluates novel ideas and hypotheses. Once a hypothesis is formulated, experiments are conducted automatically, producing results that include numerical data and visual summaries. These results are presented in tables and images, followed by an interpretation with a convincing description, culminating in a LaTeX report. In the final stage, the AI Scientist generates an automated review that refines the project and provides feedback for future open scientific discoveries. Similarly, other classic models, such as Carl [330] and Zochi [12], follow broadly analogous workflows. Notably, AgentArxiv [665] and AgentLab [666] assign distinct roles to multiple agents to simulate the collaborative nature of scientific research teams, incorporating additional peer review, academic survey, enabling semi-automatic and even full-automatic collaboration rather than relying on a single agent [478, 870, 112, 658, 53]. Despite these advancements, there remains a lack of comprehensive surveys to systematically analyze the key factors and recent developments in AI-driven research, which significantly impedes the continued progress of this field.
@@
  • Key Applications and Abundant Trending Resources: We present multidisciplinary AI4Research applications across natural sciences, applied science, and social sciences. It also identifies essential resources, open-source frameworks, public datasets, collaborative platforms, cloud-based AI services, and academic tools, that facilitate discovery management, data processing, and AI-driven research.
@@ 2 The Definition of AI4Research
Tedford [752], Society [698], [560], Farber [194], Charlin et al. [99], Charlin and Zemel [98],
@@
Figure 2: The taxonomy of AI in research (AI4Research) is categorized into five key areas. Each area is subdivided into specific tasks, underscoring the varied roles of AI in the entire research process.
@@ 2.1.3 AI for Scientific Discovery
^$\hat{\mathcal{I}}=A_{SD}(\mathcal{I})=f_{SD}(\mathcal{I}|K_{SD},R_{SD},\Phi_{SD})=f_{\text{ED}}\circ f_{\text{TA}}\circ f_{\text{NSA}}\circ f_{\text{IM}}(\mathcal{I}|K_{SD},R_{SD},\Phi_{SD})\text{,}$ (8)
@@
$\max\{\delta\}=\max\{\mathbb{E}_{\hat{\mathcal{I}}\sim A_{SD}}[\text{Novelty}(\hat{\mathcal{I}})+\text{Validity}(\hat{\mathcal{I}})+\text{Significance}(\hat{\mathcal{I}})]\},$ (9)
@@ 2.1.4 AI for Academic Writing
^$\hat{\mathcal{M}}=A_{AW}(\mathcal{M})=f_{AW}(\mathcal{M}|K_{AS},\text{Info}_{I},\Phi_{AW})=f_{\text{DWP}}\circ f_{\text{DMW}}\circ f_{\text{AWC}}(\mathcal{M}|K_{AS},\text{Info}_{I},\Phi_{AW})\text{,}$ (10)
@@
$\max\{\omega\}=\max\{\mathbb{E}_{\hat{\mathcal{M}}\sim A_{AW}}[\text{Consistency}(\hat{\mathcal{M}})+\text{Readability}(\hat{\mathcal{M}})+\text{Compliance}(\hat{\mathcal{M}})]\}\text{,}$ (11)
@@ 2.1.5 AI for Academic Peer Reviewing
^$\hat{\mathcal{R}}=A_{PR}(\mathcal{R})=f_{PR}(\mathcal{R}|P,\Phi_{PR})=f_{\text{PostP}}\circ f_{\text{InP}}\circ f_{\text{PreP}}(\mathcal{R}|\hat{\mathcal{M}},\Phi_{PR}),$ (12)
@@
$\max\{\theta\}=\max\{\mathbb{E}_{\hat{\mathcal{R}}\sim A_{PR}}[\text{Correctness}(\hat{\mathcal{R}},\hat{\mathcal{M}})+\text{Helpfulness}(\hat{\mathcal{R}},\hat{\mathcal{M}})+\text{Consistency}(\hat{\mathcal{R}},\hat{\mathcal{M}})]\}\text{,}$ (13)
@@ 3 AI for Scientific Comprehension
Scientific comprehension plays a pivotal role in advancing AI4Research, encompassing the ability to extract, understand, and synthesize information from the scientific literature. This capability not only accelerates human understanding and knowledge acquisition but also enhances the efficiency of automatic analysis, enabling more effective research processing. As shown in Figure 3, it contains two main categories: Textual Scientific Comprehension (§$\S$ 3.1) and Table & Chart Scientific Comprehension (§$\S$ 3.2).
@@ 3.1.1 Semi-Automatic Scientific Comprehension
refers to cases where a researcher’s query surpasses a language model’s knowledge base or its context-window limit [822]. The model then invokes several external tools to ensure accurate output: (1) Knowledge Retrieval Tool uses retrieval-augmented generation to inject knowledge beyond the model’s training [378, 659]. Early systems like document-centric agents [405] extract key findings, note limitations, and propose future directions. Graphusion [856] advances this with a zero-shot RAG approach: it builds scientific knowledge graphs by extracting entity triples, merging duplicates, and resolving conflicts across disciplines—without manual effort. SiGIR [135] uses self-critique feedback to guide the iterative reasoning process during knowledge-intensive multi-hop reasoning tasks. (2) Fact Checking Tool mitigates hallucinations and factual errors by applying verification modules to reduce the AI’s hallucinations [354, 278, 248, 909]. PaperQA2 [696] integrates rigorous factuality checks and matches or exceeds expert accuracy on literature-review tasks, all without unrestricted Internet access or human oversight. (3) Reasoning-Augmentation Tool addresses limited logical reasoning and computation in standalone models to deepen the AI’s theory-level comprehension [110]. For example, SciAgent [522] dynamically selects calculators and formula evaluators to deliver precise, domain-specific reasoning. Collectively, these advances show how coupling language models with specialized tools transforms scientific workflows from passive consumption into an interactive, tool-powered process that accelerates discovery while preserving rigor.
@@ 3.1.2 Full-Automatic Scientific Comprehension
More recently, LLMs have been directed to self-improve by posing clarifying questions and decomposing complex problems in a Socratic style, strengthening reasoning and conceptual understanding [622, 705, 811]. The Introspective Growth framework [829] further refines this approach, prompting smaller models to generate fundamental, open-ended questions that guide larger models toward better task comprehension. This process integrates external text retrieval to refine the understanding of technical semantics.
@@ 4 AI for Academic Survey
It is widely acknowledged that a thorough and well-conducted pre-writing survey and research phase forms the cornerstone of a successful academic article [646]. Inspired by this, AI for Academic Survey is proposed to systematically review and summarize scientific literature through the application of artificial intelligence techniques. This process plays a crucial role in ensuring that researchers and automated systems remain current with the latest advancements in their field and can efficiently identify relevant studies to inform their own work. As shown in Figure 4, it contains two main stages: Related Work Retrieval (§$\S$ 4.1) and Overview Report Generation (§$\S$ 4.2).
@@ 4.1 Related Work Retrieval
models scholarly entities (e.g., papers, authors, citations) as a graph [74, 726, 856]. Based on the types and granularity of nodes, this search method can be categorized into three types: (1) Author Relationship Graph captures the connections between researchers, enabling searches based on author relationships. For example, author-relationship graphs can effectively model collaboration networks and researcher influence [360, 597, 713]. Building on this concept, Kang et al. [359] developed a “user-recommended paper” knowledge graph that traces users’ interactions with literature, enhancing recommendation transparency and trust. (2) Paper Relationship Graph is always constructed using citation relationships between papers to construct broader paper relationships [309]. CitationSum [516] creates a citation graph linking target papers to their references with weighted relevance scores, then uses graph contrastive learning to produce abstractive summaries. (3) Entity Relationship Graph: can be constructed by modeling relationships between logical entities within papers, enabling more precise retrieval. For example, Li and Ouyang [446] propose a graph-based model that automatically identifies inter-paper entity relationships, such as contrast and support, guiding the construction of a structured Related Work section. Cross-domain graph methods are increasingly used in interdisciplinary research. Further, to address complex questions such as “Which synthesis pathways enable material X to achieve optimal conductivity?”, Ye et al. [866] systematically extract entities and their relationships from the materials science literature, thus enhancing the depth of exploration.
@@ 4.2.2 Section-level Related Work Generation
Recent studies have focused on methods to structure citations and generate cohesive connecting text for entire related work sections [789, 486, 446]. These approaches generally fall into three categories: (1) Human-Guided Generation: This approach incorporates human input, such as keywords, short abstracts, or paper groupings, to guide the generation process and maintain focus [447, 446, 533]. For instance, Gu and Hahnloser [255] and Li et al. [448] integrate user-provided or self-extracted keywords for better related-work generation. (2) Graph-Guided Generation: These methods utilize citation relationships through bibliographic graphs [226, 117, 875]. Specifically, Wang et al. [807] enhance related work generation by performing random walks on heterogeneous citation graphs. Similarly, Chen et al. [118] use a graph to link references to the paper, while Li and Ouyang [446] further prompt LLMs with inter-paper features. (3) Model-Guided Generation: In this approach, models complete the task autonomously, without additional human input [676, 790, 603]. Guo et al. [265] and Nishimura et al. [571] treat related work generation as a summarization task with structured paragraphs and novelty statements. Additionally, Pu and Demberg [605] integrate Rhetorical Structure Theory into LoRA-based fine-tuning to identify discourse relations, and Achkar et al. [4] propose a customizable multi-stage pipeline (retrieval, citation extraction, context aggregation, polishing), further enhancing the related work generation processes.
@@ 5 AI for Scientific Discovery
AI for Scientific Discovery [779, 652, 625] leverages AI to generate novel hypotheses, theories, or ideas based on existing knowledge. Its goal is to expedite the research process by automating tasks such as idea generation, novelty and significance evaluation, theoretical analysis, and experimental design. This approach not only guides new research directions but also addresses complex scientific challenges [508, 277, 284]. As shown in Figure 5, it contains five main categories: Idea Mining (§$\S$ 5.1), Novelty & Significance Assessment (§$\S$ 5.2), Theory Analysis (§$\S$ 5.3), Experiment Conduction (§$\S$ 5.4) and Full-Automatic Discovery (§$\S$ 5.5).
@@ 5.1.1 Idea Mining from Internal Knowledge
Idea mining from internal knowledge leverages the latent knowledge and generative capabilities of large language models to discover novel concepts without relying on external data [237, 684, 396]. By leveraging pretrained parameters and customized prompts, researchers can extract a variety of high-quality ideas embedded within the model [540, 104]. Earlier, Meincke et al. [541] guided LLMs toward distinct “idea spaces” by adjusting decoding temperatures and applying constraint-based prompts, effectively encouraging exploration of diverse thematic trajectories. Building on these insights, Haarmann [270] conduct an empirical study where undergraduates use an interactive tool that auto-prompted GPT-4 with business-model templates, resulting in ideas with higher novelty and feasibility, all without extensive innovation training. Additionally, Liu et al. [495] demonstrate that injecting metadata into the LLM-based ideation process and applying automated validation during selection significantly increased idea feasibility and overall quality in climate-negotiation experiments. Furthermore, Chen et al. [111] model the process of inference-time learning and reasoning as a circuit, enhancing the idea-mining ability of the model through various voltage-enhancing techniques.
@@ 5.1.2 Idea Mining from External Signal
involves supplying AI with curated academic data, such as publication metadata, citation networks, or knowledge graphs, to drive idea mining. Integrating up-to-date, domain-specific information ensures that generated hypotheses align with the latest developments in the field [483, 440, 575]. Early efforts, limited by the capacity of earlier language models, focus on predicting relationships between concepts to generate classical “A+B” ideas [288, 385]. With recent advancements in language modeling [616, 919], attention has shifted to utilizing LLMs to explore ideas from scholarly data [526, 483]. To support this, researchers have proposed various strategies for organizing literature to optimize knowledge extraction and mining. These mainly include ternary representations [793], chained structures [433], comprehensive databases [796], and knowledge graphs [79, 257, 258, 221]. Furthermore, efforts have focused on refining the knowledge injection process in idea mining. For example, Gu et al. [256] propose a framework with (1) a generalized retrieval system for cross-domain knowledge discovery and (2) a structured combinatorial process for improved idea mining.
@@
involves treating idea mining as an interactive loop that incorporates feedback from experimental or simulated environments [36, 606]. Static document mining, by contrast, often overlooks the complexities of the real world, limiting its potential for innovation. These methods enable AI systems to propose experiments, receive outcome data, and refine subsequent ideas, thereby mimicking the research cycle of design, execution, and analysis [29, 12]. In this domain, researchers primarily utilize multi-agent-based autonomous research systems, integrating idea mining agents with experiment conduction agents [666, 665, 330]. Furthermore, researchers have successfully extended idea mining to various experimental disciplines, including chemistry [518], materials science [566], biology [151], medicine [732, 249], and machine learning [320].

### v2

@@ Abstract:
[…]
@@ 1 Introduction
In recent years, the rise of artificial intelligence (AI), particularly large language models (LLMs) like DeepSeek-R1 [265], has stimulated significant research in the field of reasoning. These breakthroughs have notably improved the models’ performance across diverse areas, including mathematical reasoning, programming, and interdisciplinary knowledge [727, 751, 619, 941, 957, 111]. Some of these models have even surpassed the Turing Test [354], marking a pivotal achievement in AI development. Inspired by these, a series of works attempts to explore advanced AI systems for innovative tasks, especially in the scientific discovery of new research [869, 894, 853, 958]. Earlier, the AI Scientist [509] introduces the concept of a fully automatic AI for research system, which divides the research process into three key stages: idea mining, experiment conduction, and academic writing. Initially, the system generates and evaluates novel ideas and hypotheses. Once a hypothesis is formulated, experiments are conducted automatically, producing results that include numerical data and visual summaries. These results are presented in tables and images, followed by an interpretation with a convincing description, culminating in a LaTeX report. In the final stage, the AI Scientist generates an automated review that refines the project and provides feedback for future open scientific discoveries. Similarly, other classic models, such as Carl [332] and Zochi [12], follow broadly analogous workflows. Notably, AgentArxiv [668] and AgentLab [669] assign distinct roles to multiple agents to simulate the collaborative nature of scientific research teams, incorporating additional peer review, academic survey, enabling semi-automatic and even full-automatic collaboration rather than relying on a single agent [480, 876, 113, 661, 53]. More recently, Swanson et al. [735] extend LLMs’ scientific reach by presenting the Virtual Lab, an AI–human platform that designs nanobody binders for emerging SARS-CoV-2 variants, underscoring the promise of such collaboration for research. Despite these advancements, there remains a lack of comprehensive surveys to systematically analyze the key factors and recent developments in AI-driven research, which significantly impedes the continued progress of this field.
@@
  • Key Applications and Abundant Trending Resources: We present multidisciplinary applications of AI4Research across natural sciences, applied science, and social sciences. It also identifies essential resources, open-source frameworks, public datasets, collaborative platforms, and academic tools, that facilitate discovery management and AI-driven research.
@@ 2 The Definition of AI4Research
Tedford [755], Society [701], Nature [562], Farber [197], Charlin et al. [100], Charlin and Zemel [99],
@@
Figure 2: The taxonomy of AI for research (AI4Research) is categorized into five key areas. Each area is subdivided into specific tasks, underscoring the varied roles of AI in the entire research process.
@@ 2.1.3 AI for Scientific Discovery
$\hat{\mathcal{I}}=A_{SD}(\mathcal{I})=f_{SD}(\mathcal{I}|K_{SD},R_{SD},\Phi_{SD})=f_{\text{ED}}\circ f_{\text{TA}}\circ f_{\text{NSA}}\circ f_{\text{IM}}(\mathcal{I}|K_{SD},R_{SD},\Phi_{SD})\text{,}$ (8)
@@
[…]
@@ 2.1.4 AI for Academic Writing
$\hat{\mathcal{M}}=A_{AW}(\mathcal{M})=f_{AW}(\mathcal{M}|K_{AS},\text{Info}_{I},\Phi_{AW})=f_{\text{DWP}}\circ f_{\text{DMW}}\circ f_{\text{AWC}}(\mathcal{M}|K_{AS},\text{Info}_{I},\Phi_{AW})\text{,}$ (10)
@@
[…]
@@ 2.1.5 AI for Academic Peer Reviewing
$\hat{\mathcal{R}}=A_{PR}(\mathcal{R})=f_{PR}(\mathcal{R}|P,\Phi_{PR})=f_{\text{PostP}}\circ f_{\text{InP}}\circ f_{\text{PreP}}(\mathcal{R}|\hat{\mathcal{M}},\Phi_{PR}),$ (12)
@@
[…]
@@ 3 AI for Scientific Comprehension
Scientific comprehension plays a pivotal role in advancing AI4Research, encompassing the ability to extract, understand, and synthesize information from the scientific literature. This capability not only accelerates human understanding and knowledge acquisition but also enhances the efficiency of automatic analysis, enabling more effective research processing. As shown in Figure 3, it contains two main categories: Textual Scientific Comprehension ($\S$ 3.1) and Table & Chart Scientific Comprehension ($\S$ 3.2).
@@ 3.1.1 Semi-Automatic Scientific Comprehension
refers to cases where a researcher’s query surpasses a language model’s knowledge base or its context-window limit [827]. The model then invokes several external tools to ensure accurate output: (1) Knowledge Retrieval Tool uses retrieval-augmented generation to inject knowledge beyond the model’s training [380, 662]. Early systems like document-centric agents [407] extract key findings, note limitations, and propose future directions. Graphusion [862] advances this by building scientific knowledge graphs and extracting entity triples, resolving conflicts across disciplines without manual effort. SiGIR [137] uses self-critique feedback to guide the iterative reasoning process during knowledge-intensive multi-hop reasoning tasks. (2) Fact Checking Tool mitigates hallucinations and factual errors by applying verification modules to reduce the AI’s hallucinations [356, 280, 251, 917]. PaperQA2 [699] integrates rigorous factuality checks and matches or exceeds expert accuracy on literature-review tasks, all without unrestricted Internet access or human oversight. (3) Reasoning-Augmentation Tool addresses limited logical reasoning and computation in standalone models to deepen the AI’s theory-level comprehension [111]. For example, SciAgent [524] dynamically selects calculators and formula evaluators to deliver precise, domain-specific reasoning. Collectively, these advances show how coupling AI with specialized tools transforms scientific workflows from passive consumption into an interactive, tool-powered process that accelerates discovery while preserving rigor.
@@ 3.1.2 Full-Automatic Scientific Comprehension
More recently, LLMs have been directed to self-improve by posing clarifying questions and decomposing complex problems in a Socratic style, strengthening reasoning and conceptual understanding [625, 708, 816]. The Introspective Growth [834] further refines this approach, prompting smaller models to generate fundamental, open-ended questions that guide larger models toward better task comprehension. This process integrates external text retrieval to refine the understanding of technical semantics.
@@ 4 AI for Academic Survey
It is widely acknowledged that a thorough and well-conducted pre-writing survey and research phase forms the cornerstone of a successful academic article [649]. Inspired by this, AI for Academic Survey is proposed to systematically review and summarize scientific literature through the application of artificial intelligence techniques. This process plays a crucial role in ensuring that researchers and automated systems remain current with the latest advancements in their field and can efficiently identify relevant studies to inform their own work. As shown in Figure 4, it contains two main stages: Related Work Retrieval ($\S$ 4.1) and Overview Report Generation ($\S$ 4.2).
@@ 4.1 Related Work Retrieval
models scholarly entities (e.g., papers, authors, citations) as a graph [74, 729, 862, 716]. Based on the types and granularity of nodes, this search method can be categorized into three types: (1) Author Relationship Graph captures the connections between researchers, enabling searches based on author relationships. For example, author-relationship graphs can effectively model collaboration networks and researcher influence [362, 600]. Building on this concept, Kang et al. [361] developed a “user-recommended paper” knowledge graph that traces users’ interactions with literature, enhancing recommendation transparency and trust. (2) Paper Relationship Graph is always constructed using citation relationships between papers to construct broader paper relationships [311]. CitationSum [518] creates a citation graph linking target papers to their references with weighted relevance scores, then uses graph contrastive learning to produce abstractive summaries. (3) Entity Relationship Graph: can be constructed by modeling relationships between logical entities within papers, enabling more precise retrieval. For example, Li and Ouyang [448] propose a graph-based model that automatically identifies inter-paper entity relationships, such as contrast and support, guiding the construction of a structured Related Work section. Cross-domain graph methods are increasingly used in interdisciplinary research. Further, to address complex questions such as “Which synthesis pathways enable material X to achieve optimal conductivity?”, Ye et al. [872] systematically extract entities and their relationships from the materials science literature, thus enhancing the depth of exploration.
@@ 4.2.2 Section-level Related Work Generation
Recent studies have focused on methods to structure citations and generate cohesive connecting text for entire related work sections [793, 488, 448]. These approaches generally fall into three categories: (1) Human-Guided Generation: This approach incorporates human input, such as keywords, short abstracts, or paper groupings, to guide the generation process and maintain focus [449, 535]. For instance, Gu and Hahnloser [258] and Li et al. [450] integrate user-provided or self-extracted keywords for better related-work generation. (2) Graph-Guided Generation: These methods utilize citation relationships through bibliographic graphs [229, 118, 881]. Specifically, Wang et al. [812] enhance related work generation by performing random walks on heterogeneous citation graphs. Similarly, Chen et al. [119] use a graph to link references to the paper. (3) Model-Guided Generation: In this approach, models complete the task autonomously, without additional human input [679, 794, 606]. Guo et al. [267] and Nishimura et al. [573] treat related work generation as a summarization task with structured paragraphs and novelty statements. Additionally, Pu and Demberg [608] integrate Rhetorical Structure Theory into LoRA-based fine-tuning to identify discourse relations, and Achkar et al. [4] propose a customizable multi-stage pipeline (retrieval, citation extraction, context aggregation, polishing), further enhancing the related work generation processes.
@@ 5 AI for Scientific Discovery
AI for Scientific Discovery [783, 655, 628] leverages AI to generate novel hypotheses, theories, or ideas based on existing knowledge. Its goal is to expedite the research process by automating tasks such as idea generation, novelty and significance evaluation, theoretical analysis, and experimental design. This approach not only guides new research directions but also addresses complex scientific challenges [510, 279, 286]. As shown in Figure 5, it contains five main categories: Idea Mining ($\S$ 5.1), Novelty & Significance Assessment ($\S$ 5.2), Theory Analysis ($\S$ 5.3), Experiment Conduction ($\S$ 5.4) and Full-Automatic Discovery ($\S$ 5.5).
@@ 5.1.1 Idea Mining from Internal Knowledge
Idea mining from internal knowledge leverages the latent knowledge and generative capabilities of large language models to discover novel concepts without relying on external data [240, 687, 398]. By leveraging pretrained parameters and customized prompts, researchers can extract a variety of high-quality ideas embedded within the model [542, 105, 735]. Earlier, Meincke et al. [543] guided LLMs toward distinct “idea spaces” by adjusting decoding temperatures and applying constraint-based prompts, effectively encouraging exploration of diverse thematic trajectories. Building on these insights, Haarmann [272] conduct an empirical study where undergraduates use an interactive tool that auto-prompted GPT-4 with business-model templates, resulting in ideas with higher novelty and feasibility, all without extensive innovation training. Additionally, Liu et al. [497] demonstrate that injecting metadata into the LLM-based ideation process and applying automated validation during selection significantly increased idea feasibility and overall quality in climate-negotiation experiments. Furthermore, Chen et al. [112] model the process of inference-time learning and reasoning as a circuit, enhancing the idea-mining ability of the model through various voltage-enhancing techniques.
@@ 5.1.2 Idea Mining from External Signal
involves supplying AI with curated academic data, such as publication metadata, citation networks, or knowledge graphs, to drive idea mining. Integrating up-to-date, domain-specific information ensures that generated hypotheses align with the latest developments in the field [485, 442, 577]. Early efforts, limited by the capacity of earlier language models, focus on predicting relationships between concepts to generate classical “A+B” ideas [290, 387]. With recent advancements in language modeling [619, 927], attention has shifted to utilizing LLMs to explore ideas from scholarly data [528, 485]. To support this, researchers have proposed various strategies for organizing literature to optimize knowledge extraction and mining. These mainly include ternary knowledge representations [797], chained structures [435], comprehensive databases [800], and knowledge graphs [79, 260, 224]. Furthermore, efforts have focused on refining the knowledge injection process in idea mining. For example, Gu et al. [259] propose a framework with (1) a generalized retrieval system for cross-domain knowledge discovery and (2) a structured combinatorial process for improved idea mining.
@@
involves treating idea mining as an interactive loop that incorporates feedback from experimental or simulated environments [36, 609]. Static document mining, by contrast, often overlooks the complexities of the real world, limiting its potential for innovation. These methods enable AI systems to propose experiments, receive outcome data, and refine subsequent ideas, thereby mimicking the research cycle of design, execution, and analysis [29, 12]. In this domain, researchers primarily utilize multi-agent-based autonomous research systems, integrating idea mining agents with experiment conduction agents [669, 668, 332]. Furthermore, researchers have successfully extended idea mining to various experimental disciplines, including chemistry [520], materials science [568], biology [153], medicine [735, 252], and machine learning [322, 582].